OpenStudy (anonymous):

GH bisects CF, CF = 2y -2, and Cd = 3y -11. Find CD

OpenStudy (phi):



OpenStudy (phi):

is this the picture?

OpenStudy (anonymous):

You need to tell us more information, for example, you could draw a diagram.

OpenStudy (anonymous):

cf=2y-2 cd=3y-11 also cf=2*cd 2y-2=2(3y-11) 2y-2=6y-22 4y=20 y=5 plug value of y in cd=3y-11=3*5-11=15-11=4 so cd = 4

OpenStudy (anonymous):

How do you know that point D is the midpoint of CF?

OpenStudy (anonymous):

GH bisects CF at D so CD=DF and 2CD=CF

OpenStudy (anonymous):

The original question doesn't say GH bisects CF at D. The diagram was drawn by someone else. :(

OpenStudy (anonymous):

the question does say pl read it carefully

OpenStudy (anonymous):

GH bisects CF, CF = 2y -2, and Cd = 3y -11. Find CD. That's all the question says.